Circular No. 5941

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA CASSIOPEIAE 1993 P. Garnavich, Dominion Astrophysical Observatory, communicates:

“JHK photometry obtained with the 1.8-m Plaskett telescope indicates that the rapid brightening in the near-infrared, reported by Kidger et al. (IAUC 5936), has slowed considerably: Feb. 20.1 UT, J = 7.02, H = 4.99, K = 3.22; 21.1, 7.0, 4.8, 3.1; 22.1, 7.18, 4.93, 3.19. The uncertainties on the three nights are ±0.05, 0.1 and 0.03 mag, respectively.

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) P. W. Chodas, D. K. Yeomans and Z. Sekanina, Jet Propulsion Lab-

oratory, provide the following more detailed prediction of the impacts of the nuclei on Jupiter: nucleus 21, July 16.81 UT (corrected for light time), jovicentric latitude −43^{o}.26, earth-Jupiter-nucleus angle 98^{o}.73; 20, 17.11, −43^{o}.34, 98^{o}.49; 19, 17.27, −43^{o}.37, 98^{o}.36; 18, 17.48, −43^{o}.42, 98^{o}.20; 17, 17.61, −43^{o}.80, 98^{o}.21; 16, 18.02, −44^{o}.24, 99^{o}.42; 15, 18.30, −44^{o}.19, 98^{o}.09; 14, 18.78, −44^{o}.04, 98^{o}.13; 12, 19.42, −44^{o}.43, 97^{o}.22; 11, 19.89, −44^{o}.52, 97^{o}.58; 9, 20.41, −44^{o}.78, 96^{o}.97; 8, 20.61, −45^{o}.00, 97^{o}.47; 7, 20.80, −44^{o}.54, 96^{o}.19; 6, 21.28, −44^{o}.78, 94^{o}.39; 5, 21.61, −44^{o}.69, 95^{o}.80; 4, 21.75, −44^{o}.10, 94^{o}.80; 3, 21.88, −44^{o}.11, 94^{o}.69; 2, 22.18, −44^{o}.14, 94^{o}.46; 1, 22.32, −44^{o}.24, 94^{o}.02. The 1σ uncertainties are 0.03 day, 0^{o}.6 and 1^{o}.7, respectively.

PERIODIC COMET SCHWASSMANN-WACHMANN 2 Total visual magnitude estimates: 1993 Sept. 18.11 UT, 13.1 (B. H.

Granslo, Fjellhamar, Norway, 0.30-m reflector); Nov. 29.31, 12.3: (G. W. Kronk, Troy, IL, 0.33-m reflector); Dec. 7.21, 12.5: (Kronk); 19.04, 12.5 (K. Sárneczky, Budapest, Hungary, 0.44-m reflector); 1994 Jan. 3.84, 12.0 (L. Szentaskó, Veresegyház, Hungary, 0.33-m reflector); 11.25, 11.7 (R. Keen, Mt. Thorodin, CO, 0.32-m reflector); 19.82, 12.2 (Szentaskó); 30.15, 10.6 (C. E. Spratt, Victoria, BC, 0.20-m reflector); Feb. 2.15, 11.1 (A. Hale, Orogrande, NM, 0.20-m reflector); 8.99, 10.7 (G. Comello, Groningen, Netherlands, 0.28-m reflector); 15.14, 10.1 (J. Bortle, Stormville, NY, 0.32m reflector); 19.93, 10.2 (D. W. E. Green, Selvino, Italy, 0.25-m reflector).

PERIODIC COMET ENCKE Total magnitudes: 1993 Dec. 16.97 UT, 11.1 (Bortle); 1994 Jan. 4.11,

10.3 (Hale); 12.08, 8.8 (Keen); 20.09, 8.0 (Keen); 24.09, 7.4 (Spratt); 27.07, 7.4 (Hale); Feb. 15.77, 6.6 (D. Seargent, The Entrance, N.S.W., 25 × 100 binoculars); 16.77, 6.6 (P. Camilleri, Cobram, Victoria, 20×80 binoculars).

1994 February 24 Brian G. Marsden